OpenStudy (anonymous):

Is the line through the points R(-1,3) and S(2,-7) parallel to the graph of the line given by the equation, 10x+3y=6 No the slopes have opposite signs Yes the slopes have the same sign and value Yes the lines both decrease to the right No the lines have unequal slopes

OpenStudy (anonymous):

anything would be helpful

OpenStudy (anonymous):

First find the slope of your 2 points which is (y2- y1) / (x2-x1). Afterwards solve for y in 10x+3y=6 (so y is isolated) then find in your second equation and compare it

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ R&(-1\quad ,&3)\quad S&(2\quad ,&-7) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\)